liquid mercury (phg = 13,600 kg/m^3) is poured into a 50 cm tall beaker until the top of the mercury is 10cm below the top of the beaker.

if the air in the room is at a pressure of 1.00 atm determine the pressure of the mercury at the bottom of the beaker
1
Expert's answer
2020-12-04T07:33:10-0500
"p=p_0+\rho g(H-h)\\p=101325+(13600)(9.8)(0.5-0.1)=155000\ Pa"